# Marrying Autoregressive Transformer and Diffusion with Multi-Reference Autoregression

### v1

@@ 4 Experiments
We demonstrate through a series of experiments that TransDiff is a fast method capable of fitting image distributions. Furthermore, we validate that the novel MRAR image generation paradigm effectively enriches TransDiff’s feature representations, which enhances the diversity of generated images. Specifically, we investigate the following questions:
  • Can TransDiff beat diffusion-only and transformer-only approaches? (Table LABEL:tab:TransDiff_beats_others2)
  • Does the AR Transformer in TransDiff extract high-level semantic features, and can the diffusion decoder interpret these features to generate high-quality images? (Figure 4(a), Figure 1)
  • Is the MRAR paradigm superior to Token-Level AR and Scale-Level AR paradigms? ( Table 4.2.3, Figure 4(b))
4.1 Setup
Datasets and Metrics. We conduct comprehensive experiments on the ImageNet deng2009imagenet at $256\times 256$ and $512\times 512$ resolution, evaluating our method through quantitative metrics including Fréchet Inception Distance (FID) heusel2017gans , Inception Score (IS) salimans2016improved , Precision-Recall analysis and computational efficiency (inference time). The sampling method in the Appendix C is used in the inference stage.
Implementation Details. The Implementation Details are shown in Appendix B.
@@ 4.2 Results and Analysis
4.2.1 TransDiff beats diffusion-only and AR Transformer-only approaches
TransDiff is a unified architecture that integrates AR Transformer with diffusion model. Its core design involves generating an image representation through a single forward pass of AR Transformer, followed by a diffusion decoder to accomplish the final synthesis. In contrast, the diffusion-only approach relies entirely on a multi-step diffusion process to incrementally denoise and generate samples, while the AR Transformer-only approach sequentially constructs content through an autoregressive mechanism, generating pixels tokens one-by-one.
In the Table LABEL:tab:TransDiff_beats_others2, we compare the performance of three architectures. Under similar parameter settings, TransDiff-L with MRAR achieve the lowest FID of 1.49 and TransDiff with 1 step AR achieve a competitive remarkable speed of 0.2s/image.
Table 1: Quantitative comparison on ImageNet $256\times 256$ with existing diffusion-only and transformer-only approaches.
Type Model Params FID IS
Diffusion DiT-XL/2 675M 2.27 278.2
Diffusion MDTv2-XL 676M 1.58 314.7
AR RAR-L 461M 1.70 290.5
AR VAR-d20 600M 2.57 302.6
Hybrid TransDiff-L 1Step 683M 1.69 282.0
Hybrid TransDiff-L MRAR 683M 1.49 282.2
Model Steps FID $\downarrow$ IS $\uparrow$ Time
MAR-L 1 336.58 1.3 0.07s
MAR-L 256 1.78 296.0 1.1s
TransDiff-L 1Step 1 1.69 282.0 0.2s
TransDiff-L MRAR 4 1.49 282.2 0.8s
Diffusion Decoder vs. Diffusion Loss. In the AR transformer component, we adopt a similar architecture to MAR li2024autoregressive ; however, we employ a DitPeeblesX23 decoder to model the distribution of image samples, which fundamentally differs from the optimization objective of the MAR model based on the diffusion loss function.
To investigate the effects of the two diffusion processes, we keep the AR Transformer unchanged and perform a comparative evaluation of the differences in FID, IS, and inference time (Time) for a single image across varying inference steps between the Diffusion Loss (MAR) and the Diffusion Decoder (TransDiff). In the Table LABEL:tab:DiffdecvsDiffloss, our framework, which incorporates a diffusion decoder, demonstrates superior performance by reducing the FID from 1.78 to 1.49 and achieving faster inference. At the same time, it is important to note that our approach has the capability for one-step inference, whereas MAR does not.
State-of-the-art image generation. We evaluate TransDiff models of different sizes (Base, Large, and Huge) on the ImageNet $256\times 256$ conditional generation benchmark, and systematically compare them with state-of-the-art image generation model families. As quantitatively demonstrated in the Table 3, TransDiff achieve exceptional performance metrics characterized by extraordinarily lowest FID (e.g., 1.42) and competitive IS (e.g., 301.2). These results comprehensively validate the architectural innovation of TransDiff in simultaneously advancing both generation quality and computational efficiency, surpassing existing diffusion-based models as well as autoregressive/masked autoregressive frameworks. Additionally, these advantages hold true on the $512\times 512$ synthesis benchmark(lowest FID of 2.51)), which is provided in the Appendix D.
[…]
{tblr}
cells = c, cell11 = r=2, cell12 = r=2, cell13 = r=2, cell14 = c=2, cell16 = c=4, cell110 = r=2, vline2-4,6,10 = 1-320.05em, hline1,32 = -0.20em, hline3,6,10,14,23,26 = -0.05em, Type & Model #Params w/o CFG w/ CFG Time $\downarrow$
@@ 4.2.1 TransDiff beats diffusion-only and AR Transformer-only approaches
AR RAR-XXL DBLP:journals/corr/abs-2411-00776 1.5B - - 1.48 326.0 0.80 0.63 6.4
MAR MAR-B li2024autoregressive 208M 3.48 192.4 2.31 281.7 0.82 0.57 -
MAR MAR-L li2024autoregressive 479M 2.60 221.4 1.78 296.0 0.81 0.60 -
MAR MAR-H li2024autoregressive 943M 2.35 227.8 1.55 303.7 0.81 0.62 -
TransDiff TransDiff-B, 1 Step AR 290M 5.09 153.5 2.47 244.2 0.81 0.56 0.1
@@
TransDiff TransDiff-L, 1 Step AR 683M 3.05 185.7 1.69 282.0 0.81 0.60 0.2
@@
TransDiff TransDiff-H, 1 Step AR 1.3B 2.23 210.1 1.61 293.4 0.81 0.61 0.4
TransDiff TransDiff-H, MRAR 1.3B 2.50 200.3 1.42 301.2 0.81 0.62 1.6
@@ 4.2.2 AR Transformer’s Representation
We use the AR Transformer to extract high-level semantic features from images, which exhibit substantial representational diversity and contribute to improved performance in image generation tasks. To validate our hypothesis that more diverse image semantic features lead to higher-quality generated images, we first introduce a novel and independent metric for evaluating representational diversity. Then, we conduct comprehensive quantitative and qualitative experiments to support our conclusions.
@@
Metric for Evaluating the Diversity of Image Semantic Features. In this paper, we quantify feature diversity by computing cosine similarity matrix constructed from the $256$ L2-normalized image representations. Specifically, given the feature matrix $\mathbf{A}=[\mathbf{a}_{1};\mathbf{a}_{2};\dots;\mathbf{a}_{256}]\in\mathbb{R}^{256\times d}$, where each row $\mathbf{a}_{i}$ is an L2-normalized semantic feature vector, the cosine similarity between any two features $\mathbf{a}_{i}$ and $\mathbf{a}_{j}$ is computed as:
@@
Cosine $\text{Cosine Similarity}=\left\|\mathbf{S}\right\|_{L1}$ (15)
@@
Quantitative Analysis of Image Semantic Features. As shown in the Figure  4(a), the results demonstrate that as the number of training steps increases, the FID score consistently decreases while image diversity simultaneously improves, indicating enhanced image quality. This correlation suggests that prolonged training not only refines feature representations but also promotes greater diversity in generated samples, leading to higher-quality image generation.
Figure 4: (a):As training steps increase, the FID score consistently decreases while cosine similarity among generated samples drops, indicating improved image quality and diversity. (b): Human evaluation results among our TransDiff and other SoTA methods of Token-Level AR and Scale-Level AR.
Qualitative Analysis of Image Semantic Features
The qualitative analysis is presented in the Figure 1 . In the Middle, the results generated by the diffusion decoder using $32$, $64$, $128$, and $256$ image semantic features are shown. It can be observed that as the number of image features increases, the generated images exhibit richer details and improved visual quality, which further validates our conclusions. In addition, Bottom of the Figure  5 reveals an interesting phenomenon. When we fuse semantic features (128 features from each) from images of different classes (e.g., Macaw and Tiger), generated images exhibit characteristics of both original classes (e.g., "Tiger-striped macaw"). This observation demonstrates that features obtained through our AR Transformer indeed capture high-level semantic features of images.
4.2.3 MRAR beats Token AR and Scale-Level AR paradigms
Based on the experimental analysis in the Section 4.2.2, we observe that a higher diversity measure corresponds to a lower FID score, indicating better image quality. Therefore, we propose the MRAR paradigm (illustrated in the Figure 3 (e)) to enable the incorporation of more comprehensive feature representations.
Table 4: MRAR vs. scale AR. Compared to the baseline, MRAR achieves a higher diversity (lower diversity measure), resulting in a decrease in FID from 1.69 to 1.49. In contrast, Scale-Level AR significantly reduces the diversity (higher diversity measure) and yields a higher FID.
Model Measure $\downarrow$ FID $\downarrow$ IS $\uparrow$
TransDiff-L, 1 step AR 0.44 1.69 282.0
TransDiff-L, Scale AR 0.64 2.78 279.8
TransDiff-L, MRAR 0.39 1.49 282.2
Figure 6: Correlation between the number of references and FID score.
Human Evaluation on MRAR, Token-Level AR and Scale-Level AR.
To verify the superiority of our MRAR paradigm compared to token-level (e.g., LlamaGen-3B and RAR-XXL) and scale-level AR paradigms (e.g., VAR-d30), we conducted a human evaluation. Sixty participants ($66.7\%$ aged $24-30$, $33.3\%$ aged $30-40$; $50\%$ male, $50\%$ female; $83.3\%$ with AIGC model experience) evaluate each image using a 5-point Likert scale across three dimensions: subject diversity, background diversity, and image quality. A total of $200$ images are presented in random order to minimize bias and better capture subjective perceptions of visual diversity and quality. As shown in the Figure 4(b), participants consistently rated the MRAR paradigm higher than the token-level and scale-level AR paradigms in all three dimensions, which justifies its superiority in image generation.
Ablation Study on MRAR vs. Scale-Level AR.
To further validate the superior performance of our proposed TransDiff model with the MRAR paradigm, we fine-tuned the model using both the Scale-Level AR and MRAR paradigms based on 1-Step AR training. The experimental results in the Table 4.2.3 show that, compared to the original 1-Step AR, the MRAR paradigm achieves a higher measure, indicating greater image feature diversity, along with a lower FID score, which signifies better image generation quality. In contrast, the Scale-Level AR paradigm shows a lower measure, indicating reduced image feature diversity, and a higher FID score, indicating poorer image generation quality. This demonstrates that the MRAR paradigm, by referencing more complete and comprehensive images, is able to obtain more diverse semantic features, leading to improved generation results.
@@ 4.2.3 MRAR beats Token AR and Scale-Level AR paradigms
Properties of MRAR. To further explore the extended properties of the MRAR paradigm, we directly extend TransDiff-L into a multi-reference input framework. Through comprehensive experiments, we systematically evaluate model performance under different reference configurations (64, 16, 4, and 0), with a particular focus on analyzing the correlation between the number of references and overall image quality (i.e., fid value). The Figure 6 shows the optimal number of references (i.e., $4$).
5 Conclusion and Limitations
Conclusion: In this paper, we propose TransDiff, demonstrating through strong experiments the significant potential of marrying AR Transformers with diffusion models, without being constrained by discrete representations. Notably, our proposed MRAR paradigm showcases excellent performance, opening up a new direction in autoregressive image generation. We hope this new paradigm can inspire future research in both image and video generation.
Training Data Constraints: Primarily trained on ImageNet, the model lacks the scale and diversity of commercial proprietary datasets. This limits its ability to generate high-resolution, photorealistic images with complex semantic details, where visual quality is notably inferior to industrial models. Future work should explore larger, more diverse datasets or domain adaptation techniques to address this discrepancy.
Computational Resource Limitations: Experiments were restricted by available GPU resources, preventing investigation of larger model architectures (e.g., deeper transformers, multi-scale frameworks) or higher-resolution image generation, which could enhance quality. Hardware memory constraints were a primary factor.
Appendix A Inference Process of 1 Step AR and MRAR
The inference process of 1 Step AR is shown as the Algorithm 1:
Algorithm 1 1-Step Auto Regression inference process
Input: class token [C] and mask token $\text{Mask}_{\text{img}}$
1:  $\text{input}=\text{Concat}([\text{C},\text{Mask}_{\text{img}}])$
[…]
3:  $\text{R}_{\text{img}}=\text{Diff}(c,\epsilon)$
The inference process of MRAR is shown as the Algorithm 2:
Algorithm 2 Multi-Reference Auto Regression inference process
Input: class token [S] and mask token $\text{Mask}_{\text{img}}$
Output:the generated image $\text{R}_{\text{img}}$ corresponding to the class token [S]
1:  for $i=0$ to $n$ do
2:     if $i==0$ then
3:        $\text{input}=\text{Concat}([\text{C},\ \text{Mask}_{\text{img}}])$
4:     else
5:        $\text{input}=\text{Concat}([\text{C},\ \text{Mask}_{\text{img}},\ x_{\text{img}_{0}},...,x_{\text{img}_{i-1}}])$
6:     end if
[…]
8:     $x_{\text{img}_{i}}=c_{\text{img}_{i}}$
9:  end for
10:  $c=\text{Concat}([c_{\text{img}_{0}},c_{\text{img}_{1}},c_{\text{img}_{2}},...,c_{\text{img}_{n}}])$
11:  $\text{output}=\text{Diff}(c,\epsilon)=\text{Concat}([\text{o}_{\text{img}_{0}},\text{o}_{\text{img}_{1}},\text{o}_{\text{img}_{2}},...,\text{o}_{\text{img}_{n}}])$
12:  $\text{R}_{\text{img}}=\text{o}_{\text{img}_{n}}$
Appendix B Implementation Details.
TransDiff with 1 Step AR is trained with model sizes ranging from $290$M to $1.3$B parameters, on $32\times 8$ Nvidia A800 GPU machines with a batch size of $2048$ for $800$ epochs, using the AdamW optimizer [29] with a learning rate of $48.0e-$ 4 and a weight decay of $0.02$, along with the exponential moving average (EMA) [19] strategy. And after fine-tuning with the MRAR paradigm for 40 epochs using a separate learning rate of $55.0e-$ 5, we obtain TransDiff with MRAR.
Appendix C The Euler–Maruyama Method

### v3

@@ 4 Experiments
We demonstrate through a series of experiments that TransDiff is a efficient method capable of fitting image distributions. We validate that the novel MRAR image generation paradigm effectively enriches TransDiff’s feature representations, which enhances the diversity of generated images. Specifically, we investigate the following questions:
(I) Can TransDiff beat diffusion-only and transformer-only approaches? (Table 2)
(II) Does the AR Transformer in TransDiff extract high-level semantic features, and can the diffusion decoder interpret these features to generate high-quality images? (Figure 5, Figure 1)
(III) Is the MRAR paradigm superior to Token-Level AR and Scale-Level AR paradigms? ( Table 4, Figure 6)
4.1 Experimental Setup
Datasets and Metrics We conduct comprehensive experiments on the ImageNet deng2009imagenet at $256\times 256$ and $512\times 512$ resolution, evaluating our method through quantitative metrics including Fréchet Inception Distance (FID) heusel2017gans , Inception Score (IS) salimans2016improved , Precision-Recall analysis and computational efficiency (inference time). The Euler–Maruyama method, presented in the Appendix, is employed during the inference stage. The implementation details are shown in Appendix.
@@ 4.2 Results and Analysis
4.2.1 Comparison with diffusion-only and AR Transformer-only methods
In the Table 2, we compare the performance of three architectures among ours, diffusion only, and the AR only methods. Under similar parameter settings, TransDiff-L with MRAR achieve the lowest FID of 1.49 and TransDiff with 1-step AR achieve a competitive remarkable speed of 0.2s/image.
① Diffusion Decoder vs. Diffusion Loss. We assess the effectiveness of our usage of diffusion decoder against the the diffusion loss in MAR li2024autoregressive . Specifically, we keep the AR Transformer unchanged and perform a comparative evaluation of the differences in FID, IS, and inference time (Time) for a single image across varying inference steps between the diffusion loss (MAR) and the diffusion decoder (TransDiff). In the Table 3, our method, which incorporates a diffusion decoder, demonstrates superior performance by reducing the FID from 1.78 to 1.49 and achieving faster inference. At the same time, it is important to note that our approach has the capability for one-step inference, whereas MAR which relies on diffusion loss does not.
② Comparison between TarnsDiff and MAR. As in Figure4, it is obvious that MAR’s encoder fails to extract the semantic features while our TransDiff’s encoder exactly learn the semantic within the input images. For example, we can fuse semantic features from Icecream and Pug(128 features from each), generate ‘Pug looked icecream’.
Type Model #Params w/o CFG w/ CFG Time $\downarrow$
@@ 4.2.1 Comparison with diffusion-only and AR Transformer-only methods
MAR MAR-B li2024autoregressive 208M 3.48 192.4 2.31 281.7 0.82 0.57 0.5
MAR MAR-L li2024autoregressive 479M 2.60 221.4 1.78 296.0 0.81 0.60 1.1
MAR MAR-H li2024autoregressive 943M 2.35 227.8 1.55 303.7 0.81 0.62 2.4
TransDiff TransDiff-B, 1-Step AR 290M 5.09 153.5 2.47 244.2 0.81 0.56 0.1
@@
TransDiff TransDiff-L, 1-Step AR 683M 3.05 185.7 1.69 282.0 0.81 0.60 0.2
@@
TransDiff TransDiff-H, 1-Step AR 1.3B 2.23 210.1 1.61 293.4 0.81 0.61 0.4
TransDiff TransDiff-H, MRAR 1.3B 2.50 220.3 1.42 301.2 0.81 0.62 1.6
[…]
Type Model Params FID IS
Diffusion DiT-XL/2 675M 2.27 278.2
Diffusion DiffT 676M 1.73 276.5
AR RAR-L 461M 1.70 290.5
AR VAR-d20 600M 2.57 302.6
Hybrid TransDiff-L 1-Step AR 683M 1.69 282.0
Hybrid TransDiff-L MRAR 683M 1.49 282.2
Table 2: Comparison on ImageNet $256\times 256$ with existing diffusion-only and transformer-only approaches.
Model Steps FID $\downarrow$ IS $\uparrow$ Time
MAR-L 1 336.58 1.3 -
MAR-L 256 1.78 296.0 1.1s
TransDiff-L 1-Step AR 1 1.69 282.0 0.2s
TransDiff-L MRAR 4 1.49 282.2 0.8s
③ Comparison with SoTA Methods. We evaluate TransDiff models of different sizes (Base, Large, and Huge) on the ImageNet $256\times 256$ conditional generation benchmark, and systematically compare them with state-of-the-art image generation model families. As quantitatively demonstrated in the Table 1, TransDiff achieve exceptional performance metrics characterized by extraordinarily lowest FID (e.g., 1.42) and competitive IS (e.g., 301.2). These results comprehensively validate the architectural innovation of TransDiff in simultaneously advancing both generation quality and computational efficiency, surpassing existing diffusion-based models as well as autoregressive/masked autoregressive frameworks. Additionally, these advantages hold true on the $512\times 512$ synthesis benchmark (lowest FID of 2.51), which is provided in the Appendix.
@@ 4.2.2 AR Transformer’s Representation
To validate our hypothesis that more diverse image semantic features lead to higher-quality generated images, we first introduce a novel,independent metric for evaluating representational diversity by quantifying the L2 norm of cosine similarity matrix constructed from the 256 semantic representations for a single image. More information is shown in Appendix.
① Quantitative Analysis of Image Semantic Features. As in the Figure 5, the results demonstrate that as the number of training steps increases, the FID score consistently decreases while image diversity simultaneously improves, indicating enhanced image quality. This correlation suggests that prolonged training not only refines feature representations but also promotes greater diversity in generated samples, leading to higher-quality image generation.
Figure 5: The FID score consistently decreases while Diversity Metric among generated samples drops, indicating improved image quality and diversity.
② Qualitative Analysis of Image Semantic Features. The qualitative analysis is presented in the Figure 1 . In the Middle, the results generated by the diffusion decoder using $32$, $64$, $128$, and $256$ image semantic features are shown. It can be observed that as the number of image features increases, the generated images exhibit richer details and improved visual quality, which further validates our conclusions. In addition, bottom of the Figure 1 and Figure 4 reveals an interesting phenomenon. When we fuse semantic features (128 features from each) from images of different classes (e.g., Macaw and Tiger), generated images exhibit characteristics of both original classes (e.g., "Tiger-striped macaw"). This observation demonstrates that features obtained through our AR Transformer indeed capture high-level semantic features of images.
4.2.3 MRAR beats Token AR and Scale-Level AR paradigms
Based on the experimental analysis in AR Transformer’s representation, we observe that a higher diversity measure corresponds to a lower FID score, indicating better image quality. Thus, we propose the MRAR paradigm to enable the incorporation of more comprehensive representations.
Model Measure $\downarrow$ FID $\downarrow$ IS $\uparrow$
TransDiff-L, 1-step AR 0.44 1.69 282.0
TransDiff-L, Scale AR 0.64 2.78 279.8
TransDiff-L, MRAR 0.39 1.49 282.2
Table 4: MRAR vs. scale AR. Compared to the baseline, MRAR achieves a higher diversity (lower measure), resulting in a decrease in FID from 1.69 to 1.49. In contrast, Scale-Level AR significantly reduces the diversity.
Figure 6: Human evaluation on our TransDiff and other SoTA methods of Token-Level AR and Scale-Level AR.
① Human Evaluation on MRAR, Token-Level AR and Scale-Level AR. To verify the superiority of our MRAR paradigm compared to token-level (e.g., LlamaGen-3B and RAR-XL) and scale-level AR paradigms (e.g., VAR-d24), we conducted a human evaluation. Sixty participants ($66.7\%$ aged 243024-30, $33.3\%$ aged 304030-40; $50\%$ male, $50\%$ female; $83.3\%$ with AIGC model experience) evaluate each image using a 5-point Likert scale across three dimensions: subject diversity, background diversity, and image quality. A total of $200$ images are presented in random order to minimize bias and better capture subjective perceptions of visual diversity and quality. As shown in the Figure 6, participants consistently rated the MRAR paradigm higher than the token-level and scale-level AR paradigms in all three dimensions.
② Ablation on MRAR vs. Scale-Level AR. We fine-tune the model using both the scale-level ARtian2024visual and MRAR paradigms based on 1-Step AR training. The results in Table 4 show that, compared to 1-Step AR, MRAR achieves a higher measure, indicating greater image feature diversity, and a lower FID score, signifying better image generation quality. In contrast, the scale-level AR paradigm shows a lower measure, indicating reduced image feature diversity, and a higher FID score, reflecting poorer generation quality. This demonstrates that MRAR, by referencing more complete images, captures more diverse semantic features, leading to improved results.
5 Conclusion and Limitations
Conclusion: In this paper, we proposed TransDiff, demonstrating through strong experiments the significant potential of marrying AR Transformers with diffusion models, without being constrained by discrete representations. Notably, our proposed MRAR paradigm showcases excellent performance, opening up a new direction in autoregressive image generation. We hope this new paradigm can inspire future research in both image and video generation.
Training Data Constraints: Primarily trained on ImageNet, the model lacks the scale and diversity of commercial proprietary datasets. This limits its ability to generate high-resolution, photorealistic images with complex semantic details, where visual quality is notably inferior to industrial models. Future work should explore larger, more diverse datasets to address this discrepancy.
Appendix A Appendix
Appendix B Inference Process of 1-Step AR and MRAR
The inference process of 1-Step AR is shown as the Algorithm 1:
Algorithm 1 1-Step Autoregression inference process
Input: class token [C] and mask token Mask
[…]
The inference process of MRAR is shown as the Algorithm 2:
Algorithm 2 Multi-Reference Autoregression inference process
Input: class token [C] and mask token Mask
1:for $i=0$ to $n$ do
2:  if $i==0$ then
3:   $\text{input}=\text{Concat}([\text{C},\ \text{Mask}])$
4:  else
5:   $\text{input}=\text{Concat}([\text{C},\ \text{Mask},\ o_{\text{img}_{0}},...,o_{\text{img}_{i-1}}])$
6:  end if
7:  $c_{\text{img}_{i}}=\text{AR Transformer(input)}$
8:  $c=\text{Concat}([c_{\text{img}_{0}},c_{\text{img}_{1}},...,c_{\text{img}_{i}}])$
9:  $o_{\text{img}_{i}}=\text{Diff}(c,\epsilon)$
10:end for
[…]
Appendix C Implementation Details
TransDiff with 1-Step AR is trained with model sizes ranging from 290290M to 1.31.3B parameters, on $32\times 8$ Nvidia A800 GPU machines with a batch size of $2048$ for $800$ epochs, using the AdamW optimizer [29] with a learning rate of 8.0e48.0e-4 and a weight decay of $0.02$, along with the exponential moving average (EMA) [19] strategy. And after fine-tuning with the MRAR paradigm for 40 epochs using a separate learning rate of 5.0e55.0e-5, we obtain TransDiff with MRAR.
Appendix D Metric for Evaluating the Diversity of Image Semantic Features
@@
In this paper, we quantify feature diversity by computing cosine similarity matrix constructed from the $256$ L2-normalized image representations. Specifically, given the feature matrix $\mathbf{A}=[\mathbf{a}_{1};\mathbf{a}_{2};\dots;\mathbf{a}_{256}]\in\mathbb{R}^{256\times d}$, where each row $\mathbf{a}_{i}$ is an L2-normalized semantic feature vector, the cosine similarity between any two features $\mathbf{a}_{i}$ and $\mathbf{a}_{j}$ is computed as:
@@
$\text{Cosine Similarity}=\left\|\mathbf{S}\right\|_{L1}$ (15)
@@
Appendix E Properties of MRAR
@@
We extend TransDiff-L into a multi-reference input framework to explore the extended properties of MRAR. We systematically evaluate model performance under different reference configurations (64, 16, 4, and 0), with a particular focus on analyzing the correlation between the number of references and overall image quality (i.e., fid value). The Figure 7 shows the optimal number of references (i.e., $4$).
Appendix F The Euler–Maruyama Method